OpenStudy (anonymous):

Which one is a rule for the nth term of the sequence 3, 15, 75, 375...?

OpenStudy (solomonzelman):

\(\large\color{darkblue}{ \rm a_1-d(n-1) =a_{n}}\) ` arithmetic ` \(\large\color{darkblue}{ \rm a_1\times r^{(n-1)} =a_{n}}\) ` geometric `

OpenStudy (solomonzelman):

your sequence is geometric with the common ratio , r, of 5 and the first term, a1 of 3